OpenStudy (anonymous):

Can you Help with 4 multiple choice questions? (Will medal and fan)

OpenStudy (anonymous):

which one we doing?

OpenStudy (anonymous):

or do you have to do all?

OpenStudy (anonymous):

@satellite73 I have to complete all

OpenStudy (anonymous):

Someone Said that the first one was D but when I asked him why he couldn't tell me.

OpenStudy (anonymous):

what grade is this?

OpenStudy (anonymous):

10th

OpenStudy (anonymous):

i think i can help more with the second on

OpenStudy (anonymous):

do you know how to factorize?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ok, that one you have to factorize

OpenStudy (anonymous):

Would it be 1/3 or -5/3 ?

OpenStudy (anonymous):

-5/2 I meant. Sorry.

OpenStudy (lynfran):

i got b for the 2nd 1 6x^2+13x-5=0 6x^2+15x-2x-5=0 3x(2x+5)-1(2x+5)=0 2x+5=0 , 3x-1=0 x=-5/2 , x=1/3

OpenStudy (anonymous):

you guys are right!

OpenStudy (anonymous):

Yeah I did as well. Thats what I had chose above. Since you got the same answer as I did that means B is the solution. Can you help with the 1st, 3rd, and 4th one now?

OpenStudy (lynfran):

x^3+64=0 (x+4)(x^2-4x+16)=0 real solution =-4 so from that we can conclude that D is correct

OpenStudy (anonymous):

Ok cool I was confused for that one thank you. I'll solve 3 and 4 right now and then can you check my answer?

OpenStudy (lynfran):

ok

OpenStudy (lynfran):

for #3 u shd get 4 real solutions

OpenStudy (anonymous):

I think 3 is B

OpenStudy (lynfran):

correct

OpenStudy (anonymous):

I think 4 is C?

OpenStudy (lynfran):

ur right!!